# Collective order boundedness of sets of operators between ordered vector spaces

### v3

@@ Abstract
It is proved that: each collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded. Several applications to commutative operator semigroups on ordered vector spaces are given.
@@ 1 Introduction
Collectively compact sets of operators between normed spaces were introduced by P. M. Anselone and T. W. Palmer [5]. Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 8, 9, 10]. The present paper generalizes previous results of [1, 6, 8, 16] to collectively qualified set of operators in the ordered vector space setting. The results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator on an Archimedean ordered vector space with a generating cone is order bounded; 2) each order to norm continuous operator on an ordered Banach space with a closed generating cone is bounded.
We shall abbreviate a normed space (ordered vector space, ordered normed space, ordered Banach space, vector lattice, normed lattice, Banach lattice) by NS (OVS, ONS, OBS, VL, NL, BL). Throughout the text, vector spaces are real, operators are linear, a symbol ${\cal L}(X,Y)$ stands for the space of operators from $X$ to $Y$, and a symbol $x_{\alpha}\downarrow 0$ for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
Quite often, working with collective properties of a set of operators in OVSs involves order convergent nets. Let $X$ be an OVS. A net $(x_{\alpha})$ in $X$ is
@@
    order convergent to $x\in X$ (briefly, o-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) whenever there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is an $\alpha_{\beta}$ satisfying $-g_{\beta}\leq x_{\alpha}-x\leq g_{\beta}$ for all $\alpha\geq\alpha_{\beta}$.
@@
    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
We shall use the following OVS version of [8, Definition 2.1]. Let $X$ be an OVS, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$. We say that ${\cal B}$
@@
    collective o-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-o}}0$) if there exists a net $g_{\beta}\downarrow 0$ such that, for each $\beta$, there is an $\alpha_{\beta}$ satisfying ±$\pm x^{b}_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$ and $b\in B$.
@@
    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
@@
Let ${\cal T}\subseteq{\cal L}(X,Y)$.
When $X$ and $Y$ are OVSs,
@@
    ${\cal T}$ is collectively order bounded (briefly, ${\cal T}\in{\bf L}_{ob}(X,Y)$) if, for every order interval $[a,b]$ in $X$ there exists an order interval $[u,v]$ in $Y$ such that ${\cal T}[a,b]:=\bigcup\limits_{T\in\mathcal{T}}{T}[a,b]\subseteq[u,v]$.
@@
    ${\cal T}$ is collectively o-continuous (briefly, ${\cal T}\in\text{\bf L}_{oc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-o}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$;
@@
    ${\cal T}$ is collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
When $X$ is an OVS and $Y$ is an ONS,
  1. -
    ${\cal T}$ is collectively almost order bounded (briefly, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$) if, for every $[a,b]\subseteq X$ and $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}[a,b]\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is a normed space and $Y$ is an ONS,
  1. -
    ${\cal T}$ is collectively semicompact (briefly, ${\cal T}\in\text{\bf L}_{sc}(X,Y)$) if, for every $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}B_{X}\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is an OVS and $Y$ is a normed space,
  1. -
    ${\cal T}$ is collectively order to norm bounded (briefly, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$) if the set ${\cal T}[a,b]$ is norm bounded for every $[a,b]\subseteq X$.
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded. Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets. Finally, we prove Theorem 2.3 saying that under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded. It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded. Section 3 is devoted to applications of obtained results to operator semigroups on OVSs.
For the terminology and notations not explained in the text, we refer to [2, 3, 7, 14, 15].
2 Main results
We are going to prove the following results which generalize the results obtained in [6, 8, 16] in the vector lattice setting to the OVS setting.
  1. -
    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded.
  1. -
    Every collectively order to norm bounded set of operators from an OBS with a closed generating cone to a NS is norm bounded
  1. -
    Collectively r-continuous sets of operators between OVS with generating cones agree with collectively order bounded sets.
2.1
We begin with generalization of the first part of [8, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
@@ 2 Main results
If $X$ is an Archimedean OVS with a generating cone. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$ for each OVS $Y$.
@@
In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] as follows.
@@
Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $T$ is order bounded.
The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Theorem 2.1 and Corollary 2.1 even in the VL setting.
@@
Let $X$ be the lexicographically ordered real plane $\mathbb{R}^{2}$, $Y$ the pointwise ordered $\mathbb{R}^{2}$, and $T:X\to Y$ the identity operator. Both $X$ and $Y$ are VLs. The operator $T$ is not order bounded since the order interval $[(0,0),(2,0)]$ of $X$ contains line $x=1$ that is not order bounded in $Y$. However, $T$ is order continuous. To see this, it is enough to show that $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ implies $T((x_{\alpha},y_{\alpha}))=(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$. Suppose $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ and let $\varepsilon>0$. There exists $\alpha_{0}$ with $x_{\alpha}=0$ for $\alpha\geq\alpha_{0}$. Now, take $\alpha_{1}\geq\alpha_{0}$ such that $0\leq\max\{x_{\alpha},y_{\alpha}\}=y_{\alpha}\leq\varepsilon$ for $\alpha\geq\alpha_{1}$. It follows that $T(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$.
2.2
Next, we generalize [6, Theorem 4.6.2], [16, Theorem 5.1], and [8, Theorem 2.2] to OVSs as follows.
@@ 1. $i)$
    If $X$ has a generating cone then $\text{\bf L}_{rc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
@@ 2. $ii)$
    If $Y$ has a generating cone then $\text{\bf L}_{ob}(X,Y)\subseteq\text{\bf L}_{rc}(X,Y)$.
The idea of the proof follows the proof of [6, Theorem 4.6.2] with an additional use of the special net $x_{(k,z)}=\frac{1}{k}z$ form [1, Theorem 2.1] like in the proof of Theorem 2.1.
@@
$i)$   Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. Since $X_{+}-X_{+}=X$, for ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$. Take $b\in X_{+}$. Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1. If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$. Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{r}}0$ in $X$.
Since ${\cal T}\in\text{\bf L}_{rc}(X,Y)$, there exist $u\in Y_{+}$ and an increasing sequence $(k_{n},z_{n})$ in ${\cal I}$ satisfying ±$\pm T\big{(}\frac{z}{k}\big{)}=\pm Tx_{(k,z)}\leq\frac{1}{n}u$ for all $T\in{\cal T}$ and $(k,z)\geq(k_{n},z_{n})$. In particular, ±$\pm T\big{(}\frac{z}{k_{1}+1}\big{)}=\pm Tx_{(k_{1}+1,z)}\leq u$ for all $T\in{\cal T}$ and $z\in[0,b]$. So, ${\cal T}[0,b]\subseteq[-(k_{1}+1)u,(k_{1}+1)u]$. It follows ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
$ii)$   Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$. Then, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $-u\leq nx_{\alpha}\leq u$ (and hence $0\leq u+nx_{\alpha}\leq 2u$) for all $\alpha\geq\alpha_{n}$. Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$, there exist $w_{1},w_{2}\in Y_{+}$ satisfying $Tu\in[-w_{1},w_{1}]$ and $T(u+nx_{\alpha})\in[-w_{2},w_{2}]$ for all $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. Then $nTx_{\alpha}\in[-(w_{1}+w_{2}),(w_{1}+w_{2})]$ (and hence ±$\pm Tx_{\alpha}\leq\frac{w_{1}+w_{2}}{n}$) for $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. It follows ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. ∎
@@
Order bounded operators between OVSs with generating cones agree with r-continuous operators.
2.3
Recall that order intervals in an OBS $Y$ with a closed cone are norm bounded iff the cone is normal (cf., [3, Theorem 2.40]). We include the following elementary collective extension of this fact .
@@
Let $X$ be an OVS and $Y$ be an OBS with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
@@
Let ${\cal T}\in\text{\bf L}_{aob}(X,Y)$ and take $[a,b]\subseteq X$. Then $\bigcup_{T\in\mathcal{T}}{T}[a,b]\subseteq[-y_{1},y_{1}]+B_{Y}$ for some $y_{1}\in Y_{+}$. Since $Y_{+}$ is normal, $[-y_{1},y_{1}]+B_{Y}$ is norm bounded by [3, Theorem 2.40]. Therefore, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. ∎
The normality of $Y_{+}$ is essential in Proposition 2.1. Indeed, the identity operator in the OBS $C^{1}[0,1]$ of continuously differentiable functions on $[0,1]$ is order bounded yet not order to norm bounded.
The following result generalizes [8, Theorem 2.1] to OBSs.
@@
Let $X$ be an OBS with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
@@
Let ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. Suppose in contrary ${\cal T}B_{X}$ is not norm bounded. The M. G. Krein – Smulian theorem (cf., [3, Theorem 2.37]) implies $\alpha B_{X}\subseteq B_{X}\cap X_{+}-B_{X}\cap X_{+}$ for some $\alpha>0$, and hence ${\cal T}(B_{X}\cap X_{+})$ is not norm bounded. Pick sequences $(x_{n})$ in $B_{X}\cap X_{+}$ and $(T_{n})$ in ${\cal T}$ with $\|T_{n}x_{n}\|\geq n^{3}$ for all $n$, and set $x:=\|\cdot\|-$$\sum\limits_{n=1}^{\infty}n^{-2}x_{n}\in X_{+}$. Since ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ then ${\cal T}[0,x]\subseteq MB_{Y}$ for some $M<\infty$. It follows from $n^{-2}x_{n}\in[0,x]$ that $\|T_{n}(n^{-2}x_{n})\|\leq M$ for every $n$. This is absurd, because $\|T_{n}(n^{-2}x_{n})\|\geq n$ for all $n$. The proof is complete. ∎
It is worth noting that even in the case of a single operator, Theorem 2.3 generalizes the well known fact that every order bounded operator between BLs is bounded. The condition $X_{+}-X_{+}=X$ is essential in Theorem 2.3.

### v4

@@ Abstract
It is proved that: each collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to an ordered vector space is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded. Several applications to commutative operator semigroups on ordered vector spaces are given.
Eduard Emelyanov1, Nazife Erkurşun-Özcan2, Svetlana Gorokhova 3
$1$ Sobolev Institute of Mathematics, Novosibirsk, Russia
$2$ Hacettepe University, Ankara, Turkey
$3$ Southern Mathematical Institute, Vladikavkaz, Russia
@@ 1 Introduction
Collectively compact sets of operators between normed spaces were introduced by Anselone and Palmer [5]. Recently, collectively qualified sets of operators were investigated in [4, 10, 11, 12, 14]. In the present paper, we generalize some results of [1, 8, 10, 18] to collectively qualified sets in the ordered vector space setting. Results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator on an Archimedean ordered vector space with a generating cone is order bounded; 2) each order to norm bounded operator on an ordered Banach space with a closed generating cone is continuous.
Throughout the text, vector spaces are real, operators are linear, ${\cal L}(X,Y)$ stands for the space of operators from a vector space $X$ to a vector space $Y$, and $x_{\alpha}\downarrow 0$ stands for a decreasing net in an ordered vector space such that $\inf\limits_{\alpha}x_{\alpha}=0$.
Using of order convergence is common in study of collective properties in ordered vector spaces. A net $(x_{\alpha})$ in an ordered vector space $X$ is
@@
    order convergent to $x\in X$ (briefly, o-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) whenever there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $-g_{\beta}\leq x_{\alpha}-x\leq g_{\beta}$ for all $\alpha\geq\alpha_{\beta}$.
@@
    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
We shall use the following version of [10, Definition 1.2]. Let $X$ be an ordered vector space, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$. We say that ${\cal B}$
@@
    collective o-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-o}}0$) if there exists a net $g_{\beta}\downarrow 0$ such that, for each $\beta$, there is $\alpha_{\beta}$ satisfying ±$\pm x^{b}_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$ and $b\in B$.
@@
    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-ru}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
@@
(cf., [10, Definition 2.1]) Let ${\cal T}\subseteq{\cal L}(X,Y)$.
When $X$ and $Y$ are ordered vector spaces, the set ${\cal T}$ is
  1. $a)$
    collectively order bounded (briefly, ${\cal T}\in{\bf L}_{ob}(X,Y)$) if, for every order interval $[a,b]$ in $X$ there exists an order interval $[u,v]$ in $Y$ such that ${\cal T}[a,b]:=\bigcup\limits_{T\in\mathcal{T}}{T}[a,b]\subseteq[u,v]$.
  2. $b)$
    collectively o-continuous (briefly, ${\cal T}\in\text{\bf L}_{oc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-o}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$;
  3. $c)$
    collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-ru}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$;
When $X$ is an ordered vector space and $Y$ an ordered normed space, the set ${\cal T}$ is
  1. $d)$
    collectively almost order bounded (briefly, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$) if, for every $[a,b]\subseteq X$ and $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}[a,b]\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is an ordered vector space and $Y$ is a normed space, the set ${\cal T}$ is
  1. $e)$
    collectively order to norm bounded (briefly, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$) if the set ${\cal T}[a,b]$ is norm bounded for every $[a,b]\subseteq X$.
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to an ordered vector space is collectively order bounded. Theorem 2.4 describes relationship between collectively ru-continuous and collectively order bounded sets, and Theorem 2.8 says that, under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded. As a consequence, Corollary 2.9 says that each order to norm bounded operator from an ordered Banach space with a closed generating cone to a normed space is continuous. Proposition 2.10 tell us that every order bounded operator between ordered Banach spaces with closed generating cones is continuous. Section 3 is devoted to applications of obtained results to operator semigroups on ordered vector spaces.
For the terminology and notations that are not explained in the text, we refer to [2, 3, 9, 17].
2 Main results
We are going to prove the following results which generalize the corresponding ones obtained in [1, 10, 18] to the setting of ordered vector spaces.
@@ 2 Main results
    Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to any ordered vector space is collectively order bounded.
@@
    Theorem 2.4 saying that collectively ru-continuous sets of operators between ordered vector spaces with generating cones agree with collectively order bounded sets.
@@
    Theorem 2.8 saying that every collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space norm bounded
Let us begin with the following generalization of the first part of [10, Theorem 2.2] which is a collective version of the result by Abramovich and Sirotkin. The proof is based on ideas in the proof of [1, Theorem 2.1].
@@
Let $X$ be an Archimedean ordered vector space with a generating cone. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$ for every ordered vector space $Y$.
@@
In particular, this extends the Abramovich–Sirotkin theorem as follows.
@@
Each order continuous operator from an Archimedean ordered vector space with generating cone to any ordered vector space is order bounded.
Note that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Theorem 2.1 and in Corollary 2.2 even in the vector lattice setting.
@@
Let $X$ be the lexicographically ordered real plane $\mathbb{R}^{2}$, $Y$ the coordinate-wise ordered $\mathbb{R}^{2}$. Both $X$ and $Y$ are vector lattices. The identity operator $T:X\to Y$ is not order bounded since the order interval $[(0,0),(2,0)]$ of $X$ contains line $x=1$ that is not order bounded in $Y$. However, $T$ is order continuous. To see this, it is enough to show that $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ implies $T((x_{\alpha},y_{\alpha}))=(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$. Suppose $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ and let $\varepsilon>0$. There exists $\alpha_{0}$ with $x_{\alpha}=0$ for $\alpha\geq\alpha_{0}$. Now, take $\alpha_{1}\geq\alpha_{0}$ such that $0\leq\max\{x_{\alpha},y_{\alpha}\}=y_{\alpha}\leq\varepsilon$ for $\alpha\geq\alpha_{1}$. It follows that $T(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$.
An inspection of the proof of [1, Theorem 2.1] shows that Abramovich and Sirotkin proved already that each ru-continuous operator between vector lattices $E$ and $F$ is order bounded. Conversely, each order bounded operator $T:E\to F$ is ru-continuous since, for every $u\in E_{+}$ there exists $w\in F$ such that $T[-u,u]\subseteq[-w,w]$, and hence $|x_{\alpha}|\leq\frac{1}{n}u$ implies $|Tx_{\alpha}|\leq\frac{1}{n}w$. This assertion was independently proved by Taylor and Troitsky in [18, Theorem 5.1]. Note that this fact can also be derived from the nonstandard criteria of order boundedness of operators [8, Theorem 1.7.2] obtained by the first author. The following theorem generalizes this fact to the setting of ordered vector space.
@@ 1. $i)$
    If $X$ has a generating cone then $\text{\bf L}_{ruc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
@@ 2. $ii)$
    If $Y$ has a generating cone then $\text{\bf L}_{ob}(X,Y)\subseteq\text{\bf L}_{ruc}(X,Y)$.
@@
$i)$   Let ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$. Since $X_{+}-X_{+}=X$, to show ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$. Take $b\in X_{+}$. Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1. If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$. Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
Since ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$, there exist $u\in Y_{+}$ and an increasing sequence $(k_{n},z_{n})$ in ${\cal I}$ satisfying ±$\pm T\big{(}\frac{z}{k}\big{)}=\pm Tx_{(k,z)}\leq\frac{1}{n}u$ for all $T\in{\cal T}$ and $(k,z)\geq(k_{n},z_{n})$. In particular, ±$\pm T\big{(}\frac{z}{k_{1}+1}\big{)}=\pm Tx_{(k_{1}+1,z)}\leq u$ for all $T\in{\cal T}$ and $z\in[0,b]$. So, ${\cal T}[0,b]\subseteq[-(k_{1}+1)u,(k_{1}+1)u]$. It follows ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
$ii)$   Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$. For some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $nx_{\alpha}\in[-u,u]$ for all $\alpha\geq\alpha_{n}$. Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$ then ${\cal T}[-u,u]\subseteq[-w,w]$ for some $w\in Y_{+}$, and hence $n{\cal T}x_{\alpha}\subseteq[-w,w]$ for all $\alpha\geq\alpha_{n}$. It follows ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$. ∎
@@
Order bounded operators between ordered vector spaces with generating cones agree with ru-continuous operators.
Recall that order intervals in an ordered Banach space with a closed cone are norm bounded if and only if the cone is normal (cf., [3, Theorem 2.40]). We include the following elementary collective extension of this fact.
@@
Let $X$ be an ordered vector space and $Y$ an ordered Banach space with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
@@
Let ${\cal T}\in\text{\bf L}_{aob}(X,Y)$ and $[a,b]\subseteq X$. Then ${\cal T}[a,b]\subseteq[-y,y]+B_{Y}$ for some $y\in Y_{+}$. Since $Y_{+}$ is normal, $[-y,y]+B_{Y}$ is norm bounded by [3, Theorem 2.40]. Therefore, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. ∎
The normality of positive cone in the range is essential in Proposition 2.6.
Example 2.7.
Let $X=C^{1}[0,1]$ be an ordered Banach space of continuously differentiable functions on $[0,1]$. The identity operator $I_{X}$ is norm and order bounded yet not order to norm bounded. In particular, $I_{X}\in\text{\bf L}_{aob}(X,Y)\setminus\text{\bf L}_{onb}(X,Y)$.
The following result generalizes [10, Theorem 2.1] to ordered Banach spaces.
@@
Let $X$ be an ordered Banach space with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
@@
Let ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. Suppose, in contrary, ${\cal T}B_{X}$ is not norm bounded. The Krein – Smulian theorem (cf., [3, Theorem 2.37]) implies $\alpha B_{X}\subseteq B_{X}\cap X_{+}-B_{X}\cap X_{+}$ for some $\alpha>0$, and hence ${\cal T}(B_{X}\cap X_{+})$ is not norm bounded. Pick sequences $(x_{n})$ in $B_{X}\cap X_{+}$ and $(T_{n})$ in ${\cal T}$ with $\|T_{n}x_{n}\|\geq n^{3}$ for all $n$, and set $x:=\|\cdot\|$-$\sum\limits_{n=1}^{\infty}n^{-2}x_{n}\in X_{+}$. Since ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ then ${\cal T}[0,x]\subseteq MB_{Y}$ for some $M<\infty$. It follows from $n^{-2}x_{n}\in[0,x]$ that $\|T_{n}(n^{-2}x_{n})\|\leq M$ for every $n$. This is absurd, because $\|T_{n}(n^{-2}x_{n})\|\geq n$ for all $n$. The proof is complete. ∎
Corollary 2.9.
Every order to norm bounded operator from an ordered Banach space $X$ with a closed generating cone to a normed space $Y$ is continuous.
It is worth noting that Corollary 2.9 generalizes the well known fact that every order bounded operator between Banach lattices is bounded. It is also known that every positive operator $T:X\to Y$ is continuous whenever $X$ is an ordered Banach space with a closed generating cone and $Y$ is an ordered Banach space with a closed cone, see Arendt [6, Appendix], Aliprantis and Tourky [3, Theorem 2.32], or Arendt and Nittka [7, Theorem 2.8].
Note that a positive operator need not to be order to norm bounded. Indeed, the identity operator $I_{X}$ on an ordered Banach space $X$ is order to norm bounded if and only if the positive cone $X_{+}$ is normal. Therefore, the next proposition does not follow from Corollary 2.9. Its proof below is a modification of the proof of [3, Theorem 2.32].
Proposition 2.10.
Let $X$ and $Y$ be ordered Banach spaces with closed generating cones. Then every order bounded operator $T:X\to Y$ is continuous.
Proof.
Let $T:X\to Y$ be order bounded. To establish that $T$ is continuous, it suffices to show that $T$ has a closed graph. So, assume $x_{n}\operatorname{\xrightarrow[]{\|\cdot\|}}0$ in $X$ and $Tx_{n}\operatorname{\xrightarrow[]{\|\cdot\|}}y$ in $Y$. We must show that $y=0$.
According to [3, Lemma 2.30], there exist a subsequence $(n_{k})$ such that $x_{n_{k}}\operatorname{\xrightarrow[]{ru}}0$. By Corollary 2.5, $T$ is ru-continuous, and hence $Tx_{n_{k}}\operatorname{\xrightarrow[]{ru}}0$. Thus, there exist some $u\in Y$ and a further subsequence $(n_{k_{m}})$ such that ±$\pm Tx_{n_{k_{m}}}\leq\frac{1}{m}u$, and hence $\frac{1}{m}u\pm Tx_{n_{k_{m}}}\in Y_{+}$ for each $m$. Since $Y_{+}$ is closed and $\frac{1}{m}u\pm Tx_{n_{k_{m}}}\operatorname{\xrightarrow[]{\|\cdot\|}}\pm y$, we conclude that ±$\pm y\in Y_{+}$. That is, y = 0 and the proof is finished. ∎
We have the following collective extension of Proposition 2.10.
Proposition 2.11.
Let $X$ and $Y$ be ordered Banach spaces with closed generating cones, and $X^{\sim}\neq\{0\}$. The following assertions are equivalent.
  1. $i)$
    Every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is norm bounded.
  2. $ii)$
    The cone $Y_{+}$ is normal.
Proof.
$i)\Longrightarrow ii)$ Assume that $Y_{+}$ is not normal. Then, there exists $y_{0}\in X_{+}$ such that the interval $[0,y_{0}]$ is not bounded. Pick $f_{0}\neq 0$ in $X^{\sim}$ and take $x_{0}\in X$ with $f_{0}(x_{0})=1$. Then, $\{f_{0}\otimes y\}_{y\in[0,y_{0}]}\in\text{\bf L}_{ob}(X,Y)$. However, the set $\{f_{0}\otimes y\}_{y\in[0,y_{0}]}$ is not norm bounded because of $\bigcup\limits_{y\in[0,y_{0}]}(f_{0}\otimes y)(x_{0})=[0,y_{0}]$.
$ii)\Longrightarrow i)$ Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$. Then, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$. So, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ by Proposition 2.6. It follows from Theorem 2.8 that ${\cal T}$ is norm bounded. ∎
The following example shows that the condition that $X_{+}$ is closed and generating is essential in Theorem 2.8, in Corollary 2.9, and in Proposition 2.10.